OpenStudy (anonymous):

What are the factors of x2 − 64? Prime (x − 4)(x + 16) (x − 8)(x − 8) (x + 8)(x − 8)

OpenStudy (harsimran_hs4):

use: a^2 - b^2 = (a-b)(a+b)

OpenStudy (harsimran_hs4):

@teddybearbloom did you get started?

OpenStudy (primeralph):

Try expanding each one.

OpenStudy (anonymous):

ughh no idk what a and b are

OpenStudy (harsimran_hs4):

in your ques i can clearly see a atleast , it`s x so moving along same lines can you spot b ?

OpenStudy (anonymous):

the answer is (x + 8)(x − 8)

OpenStudy (anonymous):

difference of squares \[a^2-b^2 = (a-b)(a+b)\] i suggest memorizing this

OpenStudy (anonymous):

but yes you have it right in your case, a= x and b=8